let A be a 7*5 matrix with rank(A)=2 complete dim(row space of A) , dim( column space of A) ,dim (null space of A) and (null space of A^t)


1
Expert's answer
2021-09-14T06:09:29-0400

dim(row space of A)

"dim (rowspace(A)) = rank(A) = dim (colspace(A))\\
dim (rowspace(A))+dim(null (A))=7\\
dim (rowspace(A))=7-dim(col (A))\\
dim (rowspace(A))=7-4=3\\"


dim( column space of A)

"[A]_{5*7}" gives us a transformation T with domain v of dimension T

"dim(col (A))+dim(null (A))=7\\
dim(col (A))=7-dim(col (A))\\
dim(col (A))=7-4=3\\"


dim (null space of A)

"dim(null (A))+dim(col (A))=7\\
dim(null (A))=7-dim(col (A))\\
dim(null (A))=7-5=2\\"


(null space of At)

"(null space of A^t)= (null space of A)=2"